Warren has 40 coins (all nickels, dimes, and quarters) worth $4.05. He has 7 more nickels than dimes. How many quarters does Warren have?
Answer by stanbon(48550)

Warren has 40 coins (all nickels, dimes, and quarters) worth $4.05. He has 7 more nickels than dimes. How many quarters does Warren have?
---------
Let number of dimes be "x"; their value is 10x cents.
Number of nickels is "x+7"; their value is 5(x+7)=5x+35 cents
Number of quarters is "40-x-(x+7)" = "33-2x" ; their value is 25(33-2x)
=825-50x cents
EQUATION:
value + value = value
10x+5x+35+825-50x=405 cents
-35x=-45
x=13 (number of dimes)
x+7=20 (number of nickels)
33-2x=33-26=7 (number of quarters)
Cheers,
Stan H.